# FlashBack : Efficient Retrieval-Augmented Language Modeling for Long Context Inference

### v1

@@ Abstract
Retrieval-Augmented Language Modeling (RALM) by integrating large language models (LLM) with relevant documents from an external corpus is a proven method for enabling the LLM to generate information beyond the scope of its pre-training corpus. Previous work using utilizing retrieved content by simply prepending retrieved contents to the input poses a high runtime issue, which degrades the inference efficiency of the LLMs because they fail to use the Key-Value (KV) cache efficiently. In this paper, we propose FlashBack, a modular RALM designed to improve the inference efficiency of RALM with appending context pattern while maintaining decent performance after specific fine-tuning without heavily destruct the knowledge integrity of the LLM. FlashBack appends retrieved documents at the end of the context for efficiently utilizing the KV cache instead of prepending them. Our experiment shows that the inference speed of FlashBack is up to 4×4$\times$4 × faster than the prepending method on a 7B LLM (Llama 2). Via bypassing unnecessary re-computation, it demonstrates an advancement by achieving significantly faster inference speed, and this heightened efficiency will substantially reduce inferential cost. Our code will be publicly available.
@@ 1 Introduction
However, these works are introduced with limitations. First, the off-the-shelf LLMs are not inherently trained to incorporate retrieved content, and extensive pre-training of LLMs for building RALM incurs high computational costs (Lin et al., 2023). Second, although in-context methods have been effectively applied on off-the-shelf LLMs (Ram et al., 2023; Shi et al., 2023), recent research indicates that the bottleneck of these methods is redundancy and inefficiency (Asai et al., 2024) and our work verified that prepending context pattern they used poses high computational cost in long context scenario, which tremendously increases the inference time. To be specific, retrieved content changes with every retrieval stride tokens, and the computed key-value (KV) cache of the entire context is discarded, requiring re-computation for each change. Text generation process with long context can substantially increase the inference runtime, and the attention KV cache grows even larger on LLM with hundreds of billions of parameters(Pope et al., 2022). Under these circumstances, the re-computation of the KV cache becomes a computationally intensive task, entailing significant costs for In-Context RALM(Ram et al., 2023).
In this paper, we propose FlashBack to overcome these issues we mentioned above, a method that utilize off-the-shelf LLMs efficiently without further pre-training of LLMs as shown in figure 3, to address the inefficiency issue by breaking the dependent of the input to retrieved content in the context. To be specific, instead of prepending the retrieved content directly to the input, we propose a new context pattern, Appending Context Pattern, in which the retrieved content is appended at the end of the input (see figure 2). In our method, the KV cache of each input is static during inference, and the re-computation of the KV cache takes place only in the part of the retrieved content. Although there are advantages to applying this new context pattern, it also presents an issue that degrading the performance of the RALM. Our appending pattern breaks the semantic coherence of the context, which leads to a loss in perplexity when applying it to existing RALM. To solve this issue, inspired by (Ren et al., 2023), FlashBack introduces tunable Marking Token to adapt to our appending context pattern for restoring performance, which is different from pre-training.
Notably, we use LoRA (Hu et al., 2021) to fine-tune the Marking Token while keeping both the retriever and the LLM frozen. Therefore, FlashBack is a plug-and-play approach that can cooperate with other retrieval-oriented methods aiming for efficient inference and context-model alignment. When using an LLM with 7 billion parameters (7B) as a reader, user can run the whole token-tuning process with a single 24GB-RAM GPU in BF16 precision.
@@
  • We exploit Marking Token that can be used for adapting the LLMs to unseen context pattern by slight fine-tuning with LoRA. And this alignment method with LoRA can keep the LLM intact and avoid heavily damaging the knowledge integrity of the LLM.
@@ 3.2 Context pattern
Figure 5: Time for OPT models generating a sequence of length 2048 tokens (including 128 tokens for retrieved content), given a input of 512 tokens, with retrieval stride of 16.
@@ 3.4 Marking Token and Fine-tuning Choice
Since LLMs are not aligned explicitly to our appending pattern, we use Marking Token and LoRA techniques to adapt them to the appending pattern while keeping origin model weights frozen so that the alignment is achieved without modifying the inherent ability of the LLMs. Marking Tokens are represented as two special prompt tokens denoting <MARK_L>,<MARK_R> which are added into the vocabulary of the model, and they are used for marking the boundary of the retrieved content in context (see Figure 3). Then, we fine-tune the model using Low-Rank Adaption (LoRA) (Hu et al., 2021) with Marking Tokens applied for adapting the appending pattern. LoRA is a parameter-efficient fine-tuning (PEFT) for fine-tuning a subset of parameters in the pre-trained model while targeting for obtaining comparable performance to full model fine-tuning. Recent study suggest that forgetting is inevitable in fine-tuning but PEFT method can facilitate less forgetting during fine-tuning process which does not heavily damage the knowledge integrity of the pre-trained LLM (Kalajdzievski, 2024).In FlashBack, we opt for LoRA with the aim of demonstrating how fine-tuning a relatively minimal set of model weights enables our model to adapt to the appending context pattern while keeping pre-training model weights frozen. Our experiments show that FlashBack has significantly boosted the inference speed, and it can still maintain decent or better perplexity (PPL) after fine-tuning. During fine-tuning, all parameters are frozen except for the embedding of <MARK_L>, <MARL_R>, and LoRA modules applied to attention layers.
@@ 3.6 Retriever
Table 1: Perplexity (the lower the better) of GPT2 and OPT models on a part of the validation set of WikiText-2 dataset and the test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16. A. and P. refers to the appending and prepending context pattern respectively. M.T. refer to the Marking Tokens. And N. refers to none retrieval.
4 Experiments
4.1 Run-time Improvement
@@
Figure 6: Time for Llama2-7B generating sequences with different length, given a input of 512 tokens, with retrieval stride of 16.
@@ Results
In Figure 5, we compare the appending pattern and prepending pattern in OPT models. We scale the model size from 125M to 6.7B. It is obvious that the acceleration of inference time is more effective in large models that have more layers and larger hidden sizes. In figure 6, we scale the max sequence length from 1088 to 3968 and observe significant improvement, as Llama 2 has a larger max sequence length, which is 4096 tokens. (Note: The test results for inference speed may vary depending on the hardware used; our runtime tests are conducted on a single A100-80GB GPU)
4.2 Language Modeling
Setup
We test FlashBack on Wikitext-2 (Merity et al., 2016) dataset, and three dataset: Arxiv, Freelaw and Stackexchange from the Pile (Gao et al., 2020). We report token-level average perplexity as the evaluation metric (excluding <MARK_L>,<MARK_R>). Our experiments use models of OPT (1.3B-6.7B) and GPT-2 (124M-1.5B). We compare FlashBack with the appending context pattern without fine-tuning and also the prepending context pattern. (For smaller models that are under 774M parameters, the results are in Appendix Table 6)
Results
Table 1 shows the results of different setup on the dev set. We can see that the appending context pattern applied directly on LLMs through the in-context method generates higher perplexity compared with the model using the prepending context pattern. Our result on fine-tuning with appending context patterns by using LoRA demonstrates that the model can adapt to such patterns after fine-tuning. Notably, with both the Marking Token and LoRA applied, the difference between perplexity results of two different context patterns is increasingly mitigated as the test model size grows larger.
@@ 4.4 Implementation Details
The trainable parameters for all models include two token embedding of <MARK_L> and <MARK_R>, and LoRA modules applied to all attention layers. The rank of weight matrices of LoRA is set to 16. The batch size is set to 16. The learning rates for GPT-2 models (124M-1.5B) are specified as [4e-4, 2e-4, 1e-4, 5e-5]. Correspondingly, the learning rates for OPT models (1.3B-6.7B) are specified as [1e-4, 5e-5, 5e-5]. We train for 5K additional steps, with 10% warm-up steps and a cosine learning rate schedule. We use the bf16 format for the OPT-6.7B model and set the maximum sequence length to 512 for running on a single 24GB RAM GPU.
Our code uses the Transformers library (Wolf et al., 2020) and BM25 retrieval from the Pyserini library (Lin et al., 2021). All fine-tuning experiments are conducted by using 4$\times$24GB RAM GPUs.
@@ 5 Conclusion
This paper presented the FlashBack method, enabling In-Context RALM with faster inference speed while keeping competitive model performance. We proved that the appending context pattern is more efficient than the prepending counterpart. Besides, we demonstrated that LLMs can adapt to new context patterns using Marking Tokens with adaption fine-tuning (LoRA) and its performance in generation quality is likely equivalent to the LLM using prepending counterpart as model grows larger.
@@ 6 Limitations
In our experiments, the retrieval stride is a fixed value in each test, but it can be a dynamic variable in real applications, so there might be extra improvements. In run-time comparison experiments, due to constrained computational resources, we used simulated (pseudo) data for testing; therefore, we have results for comparing inference speed. Also, It is computationally intensive to fine-tune an LLM over 7B parameters, so the performance of FlashBack in extra large models (over 7B) might require further demonstration.
Frozen LM with LoRA modules explicitly makes it easy to switch modes between LLM and RALM. By just loading LoRA modules, a frozen LLM can use the external corpus and vice versa. In some situations, knowledge stored in model parameters might be superior so that it is possible for RALM to evaluate the quality of the retrieved contents and dynamically determine whether the external corpus should be used in generating the output. Furthermore, in FlashBack, we did not scale the number of retrieved contents to a large value. It is practicable to generalize appending context patterns to more retrieved contents (for example, using the framework of REPLUG(Shi et al., 2023)), which can be one of the future work to do.
These we mentioned above can be future research for RALM.
@@ Appendix A FlashBack Algorithm
Appendix B Setup and Experiments
@@ B.1 Ablation Study
B.2 Perplexity of GPT model in small model size
@@
Table 6: Perplexity (the lower the better) of GPT2 models on a part of the validation set of WikiText-2 dataset and the test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16. A. and P. refers to the appending and prepending context pattern respectively. M.T. refers to the Marking Tokens. And N. refers to none retrieval.

### v3

@@ Abstract
Retrieval-Augmented Language Modeling (RALM) by integrating large language models (LLM) with relevant documents from an external corpus is a proven method for enabling the LLM to generate information beyond the scope of its pre-training corpus. Previous work utilizing retrieved content by simply prepending it to the input poses a high runtime issue, which degrades the inference efficiency of the LLMs because they fail to use the Key-Value (KV) cache efficiently. In this paper, we propose FlashBack, a modular RALM designed to improve the inference efficiency of RALM with appending context pattern while maintaining decent performance after fine-tuning by Low-Rank Adaption. FlashBack appends retrieved documents at the end of the context for efficiently utilizing the KV cache instead of prepending them. And we introduce Marking Token as two special prompt tokens for marking the boundary of the appending context during fine-tuning. Our experiments on testing generation quality show that FlashBack can remain decent generation quality in perplexity. And the inference speed of FlashBack is up to 4×4$\times$4 × faster than the prepending counterpart on a 7B LLM (Llama 2) in the runtime test. Via bypassing unnecessary re-computation, it demonstrates an advancement by achieving significantly faster inference speed, and this heightened efficiency will substantially reduce inferential cost.
@@ 1 Introduction
However, these works are introduced with limitations. First, the off-the-shelf LLMs are not inherently trained to incorporate retrieved content, and extensive pre-training of LLMs for building RALM incurs high computational costs (Lin et al., 2023). Second, although in-context methods have been effectively applied on off-the-shelf LLMs (Ram et al., 2023; Shi et al., 2023), recent research indicates that the bottleneck of these methods is redundancy and inefficiency (Asai et al., 2024). And our work analyze that prepending context pattern they used poses high computational cost in long context scenario, which tremendously increases the inference time (see Figure 1). To be specific, retrieved content changes with every retrieval stride tokens, and the computed key-value (KV) cache of the entire context is discarded, requiring re-computation for each change. Text generation process with long context can substantially increase the inference runtime, and the attention KV cache grows even larger on LLM with hundreds of billions of parameters(Pope et al., 2022). Under these circumstances, the re-computation of the KV cache becomes a computationally intensive task, entailing significant costs for In-Context RALM(Ram et al., 2023).
In this paper, we propose FlashBack to overcome these issues we mentioned above, a method that fine-tuning off-the-shelf LLMs efficiently through adaption without further pre-training of LLMs as shown in figure 2, and to address the inefficiency issue by breaking the dependent of the input to retrieved content in the context. To be specific, instead of prepending the retrieved content directly to the input, we propose a new context pattern, Appending Context Pattern, in which the retrieved content is appended at the end of the input (see figure 3). In our method, the KV cache of each input is static during inference, and the re-computation of the KV cache takes place only in the part of the retrieved content. Although there are advantages to applying this new context pattern, it also presents an issue that degrading the performance of the RALM. Our appending pattern breaks the semantic coherence of the context, which leads to a loss in perplexity when applying it to existing RALM. To solve this issue, inspired by (Ren et al., 2023), FlashBack introduces tunable Marking Token to adapt to our appending context pattern for restoring performance, which is different from pre-training the entire model weights.
Notably, we use LoRA (Hu et al., 2021) to fine-tune the Marking Token while keeping both the retriever and the LLM frozen. Therefore, FlashBack is a plug-and-play approach that can cooperate with other retrieval-oriented methods aiming for efficient inference and context-model alignment. When using an LLM such as OPT-6.7B as a reader, user can run the whole token-tuning process with a single 24GB-RAM GPU in BF16 precision.
@@
  • We exploit Marking Token that can be used for adapting the LLMs to unseen context pattern by fine-tuning with LoRA. And this alignment method with LoRA can keep the LLM intact and avoid heavily damaging the knowledge integrity of the LLM.
@@ 3.4 Marking Token and Fine-tuning Choice
Since LLMs are not aligned explicitly to our appending pattern, we use Marking Token and LoRA techniques to adapt them to the appending pattern while keeping origin model weights frozen so that the alignment is achieved without modifying the inherent ability of the LLMs. Marking Tokens are represented as two special prompt tokens denoting <MARK_L>,<MARK_R> which are added into the vocabulary of the model, and they are used for marking the boundary of the retrieved content in context (see Figure 2). Then, we fine-tune the model using Low-Rank Adaption (LoRA) (Hu et al., 2021) with Marking Tokens applied for adapting the appending pattern. LoRA is a parameter-efficient fine-tuning (PEFT) for fine-tuning a subset of parameters in the pre-trained model while targeting for obtaining comparable performance to full model fine-tuning. Recent study suggests that forgetting is inevitable in fine-tuning but PEFT method can facilitate less forgetting during fine-tuning process which does not heavily damage the knowledge integrity of the pre-trained LLM (Kalajdzievski, 2024). In FlashBack, we opt for LoRA with the aim of demonstrating how fine-tuning a relatively minimal set of model weights enables our model to adapt to the appending context pattern while keeping pre-training model weights frozen. Our experiments show that FlashBack has significantly boosted the inference speed, and it can still maintain competitive perplexity (PPL) after fine-tuning. During fine-tuning, all parameters are frozen except for the embedding of <MARK_L>, <MARL_R>, and LoRA modules applied to attention layers.
@@ 3.6 Retriever
4 Experiments
4.1 Run-time Improvement
Setup
Results
In Figure 5, we compare the appending pattern (in-context), prepending pattern (in-context) and FlashBack in OPT models. We scale the model size from 125M to 6.7B. It is obvious that the acceleration of inference time is more effective in large models that have more layers and larger hidden sizes. In figure 6, we scale the max sequence length from 1088 to 3968 and observe significant improvement, as Llama 2 has a larger max sequence length, which is 4096 tokens. (Note: The test results for inference speed may vary depending on the hardware used and setting; our runtime tests are conducted on a single A100-80GB GPU setting to P0 state)
Figure 5: Time for OPT models generating a sequence of length 2048 tokens (including 128 tokens for retrieved content), given a input of 512 tokens, with retrieval stride of 16. We test appending and prepending patterns through in-context method for comparing with FlashBack. This is also a part of the ablation study.
4.2 Language Modeling
Figure 6: Time for Llama2-7B generating sequences with different length, given a input of 512 tokens, with retrieval stride of 16. This figure is for showing how the runtime increases in each distinct context pattern as sequence length scales up.
@@ 4.2 Language Modeling
Table 1: Perplexity (the lower the better) of OPT models on the entire validation set of WikiText-2 dataset and partial test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16. A. and P. refers to the appending and prepending context pattern respectively. M.T. refer to the Marking Tokens. And N. refers to none retrieval.
@@
We test FlashBack on Wikitext-2 (Merity et al., 2016) dataset, and three dataset: Arxiv, Freelaw and Stackexchange from the Pile (Gao et al., 2020). We report token-level average perplexity as the evaluation metric (excluding <MARK_L>,<MARK_R>). Our experiments use models of OPT (1.3B-6.7B) and GPT-2 (124M-1.5B). We compare FlashBack with the appending context pattern without fine-tuning and also the prepending context pattern. (For smaller models that are under 774M parameters, the results are in Appendix C Table 6)
@@ Results
Table 1 shows the results of different setup on the dev set. We can see that the appending context pattern applied directly on LLMs through the in-context method generates higher perplexity compared with the model using the prepending context pattern. Our result on fine-tuning with appending context patterns by using LoRA demonstrates that the model can adapt to such patterns after fine-tuning. Notably, with both the Marking Token and LoRA applied, the difference of perplexity results between two distinct context patterns becomes progressively alleviated as the size of the test model increases. And we also test GPT-2 from 124M to 1.5B to validate this observation. (See Table 6 in Appendix D)
@@ 4.4 Implementation Details
The trainable parameters for all models include two token embedding of <MARK_L> and <MARK_R>, and LoRA modules applied to all attention layers. The rank of weight matrices of LoRA is set to 16. The batch size is set to 16. The learning rates for GPT-2 models (124M-1.5B) are specified as [4e-4, 2e-4, 1e-4, 5e-5]. Correspondingly, the learning rates for OPT models (1.3B-6.7B) are specified as [1e-4, 5e-5, 5e-5]. We train for 5K additional steps, with 10% warm-up steps and a cosine learning rate schedule. For running on a single 24GB RAM GPU, it is feasible to use the bf16 format on the OPT-6.7B model and set the maximum sequence length to 512.
In our implementation, we uses the Transformers library (Wolf et al., 2020) and BM25 retrieval from the Pyserini library (Lin et al., 2021). And all fine-tuning experiments are conducted by using 4$\times$24GB RAM GPUs.
@@ 5 Conclusion
This paper presented the FlashBack method, enabling RALM with faster inference speed while keeping competitive model performance. We proved that the appending context pattern is more efficient than the prepending counterpart. Besides, our experiments demonstrated that LLMs can adapt to new context patterns using Marking Tokens with adaption fine-tuning (LoRA). And its performance in generation quality is closely equivalent to the LLM using prepending counterpart as size of the model increases and.
@@ 6 Limitations
In our experiments, the retrieval stride is a fixed value in each test, but it can be a dynamic variable, so there might be extra improvements. In run-time comparison experiments, due to constrained computational resources, we used simulated (pseudo) data for testing. It is because we did not find an appropriate public downstream task which fulfill following requirements: 1) long context, 2) multiple retrieval, 3) generated text becomes a new part of the context. However, we do believe this type of task exists in industrial applications. Additionally, we did not test our modeling on LLMs over size of 7B parameters since they generally require extensive computational resource.
Furthermore, in FlashBack, we did not scale the number of retrieved contents to a large value. It is practicable to generalize appending context patterns to more retrieved contents (for example, using the framework of REPLUG(Shi et al., 2023)).
Appendix
@@ Appendix A FlashBack Algorithm
Appendix B Experiment Setup
@@ Appendix C Ablation Study
Appendix D Perplexity results of GPT2 model
@@
GPT2-774M N. 19.11 10.69 11.29 9.74
P. 16.71 10.58 11.08 9.59
P. 11.74 8.56 8.03 6.45
P. 11.69 8.52 8.03 6.43
A. 56.15 42.44 49.25 39.61
A. 14.40 16.83 16.35 11.99
A. 12.13 8.95 8.47 6.78
GPT2-1.5B N. 17.32 9.87 10.61 9.03
P. 15.36 9.81 10.49 9.01
P. 10.84 8.07 7.45 6.01
P. 10.83 8.04 7.43 5.98
A. 53.30 41.72 49.25 40.25
A. 13.75 16.48 15.20 11.50
A. 11.32 8.57 8.01 6.35
Table 6: Perplexity (the lower the better) of OPT models on the entire validation set of WikiText-2 dataset and partial test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16. A. and P. refers to the appending and prepending context pattern respectively. M.T. refer to the Marking Tokens. And N. refers to none retrieval.